(2) The balance in your saving account on 01/01/2013 was $12,000. On 01/01/2016 you check the balance again, and it is S12,894. The bank computes the interest rate compounded monthly What is the interest rate per month? (Hint 1: First, find how many months there are between two dates. Note the question is asking the interest rate per month, not per year. Solve for the r using the above formula. Hint 2: The formula FV C) is equivalent to 1 +r) 1 . 0.10% 2. 0.1 5% 3.0.20% 4.0.25% 5. 0.30%
Answer #1

Using the formula,

FV = C(1+r)T

since it is compounded monthly, r becomes r/12 and T becomes nT i.e 3T since 3 years

FV = C(1+r/12)3T

12894 = 12000(1+r/12)36

r/12 = (12894/12000)1/36 - 1 = 0.2%

r/12 = monthly interest rate = 0.2%